OpenStudy (anonymous):

A jar contains 26 marbles. It has 10 red, 8 black, and 8 green marbles. Two marbles are drawn; the first is not returned before the second one is drawn. What is the probability that both marbles are green? P(both green) = 8/13 P(both green) = 28/325 P(both green) = 32/325 P(both green) = 14/169

OpenStudy (anonymous):

I think the answer is 14/169. 8/13 times 7/26 equals 14/169

OpenStudy (anonymous):

How did you get there?

OpenStudy (anonymous):

Like how did you 8/13 and 7/26?

OpenStudy (anonymous):

There is 8 marbles out of 26 that are green. After picking out a marble from those 8 marbles, there are only 7 marbles that are green. Thats how I got there

OpenStudy (anonymous):

Ok. I have a few more, would you mind helping me?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

But I'm a bit slow since I'm helping 2 people including you.

OpenStudy (anonymous):

By accident, six burned-out bulbs have been mixed in with 26 good ones. Ken is replacing old bulbs in his house. If he selects two bulbs at random from the box of 32, what is the probability they both work? P(both work) = 325/512 P(both work) = 325/496 P(both work) = 169/256 P(both work) = 169/248

OpenStudy (anonymous):

Okay that's fine!

OpenStudy (anonymous):

I think its A

OpenStudy (anonymous):

can you explain?

OpenStudy (anonymous):

There are 26 bulbs out of 32 bulbs that work. If he takes out one, there is 25 left. So I multiplied 13/16 (simplified version of 26/32) and 25/32. Oh wait let me revise

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

The answer is B. There are 26 bulbs out of 32 bulbs that work. If he takes out one, there are 25 left out of 31 bulbs. I multiplied 13/16 and 25/31

OpenStudy (anonymous):

Ok great! thank you!

OpenStudy (anonymous):

No problem. Can you fan+medal

OpenStudy (anonymous):

Yes!